OpenStudy (anonymous):

find angle A if side a=34 side b=28 and side c=15 need help bad please help me

OpenStudy (anonymous):

 do you know the Law of Cosines?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

there are actually 3 but they're all basically the same... since you want Angle A, i'll write it out for you in it's most common form: a^2 = b^2 + c^2 -2abcosA

OpenStudy (anonymous):

using this formula, solve for cosA... what do you have?

OpenStudy (anonymous):

34 is cos a

OpenStudy (anonymous):

forget about the lengths given for now... solve for cosA in terms of a, b, c.

OpenStudy (anonymous):

ok so if I have a 34 b 28 and c 15 would i have to times them all by their number again

OpenStudy (anonymous):

this gives you cosA. Do you know how to find A from here?

OpenStudy (anonymous):

ya so if I add bsquared and c swuared I would get 666

OpenStudy (anonymous):

b^2 + c^2 = 666 ?

OpenStudy (anonymous):

ok I have no idea how to do it can you just do it really quick so I can learn from you how to do it please that would help a lot

OpenStudy (anonymous):

hang on, there's a problem here...

OpenStudy (anonymous):

 ok, that should be it.